If the area of a square is 4 meter square then what is the length of its diagonals?

Answer by TimothyLamb(4379)

a = LL = 4
L = 2 m
---
d = sqrt( LL + LL )
d = sqrt( 2LL )
d = sqrt( 8 )
---
d = 2.8284271 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php